# Scattering Theory of Chiral Edge Modes in Topological Magnon Insulators

### v1

@@ 1 Numerical Extraction of Scattering Phases
2 Effective Theory on the Edge
In the main text we mention the possibility of constructing an effective one-dimensional edge theory for the scattering magnons from the knowledge of the scattering phase. In this section we present some further details for clarity. Since chiral edge modes remain confined in the lowest band, we can effectively project the Hamiltonian within this subspace, and introduce the effective two-particle Hamiltonian
$\mathcal{H}_{\text{2-eff}}=\varepsilon(k_{1})+\varepsilon(k_{2})+U_{\text{2-eff}}(k_{1},k_{2})\,,$ (S3)
where $U_{\text{2-eff}}(k_{1},k_{2})$ captures the effective interactions among the two chiral magnons of energies $\varepsilon(k_{1,2})$. The effective one-dimensional theory must lead to the same scattering events of the full two-dimensional scattering; this condition uniquely fixes $U_{\text{2-eff}}(k_{1},k_{2})$ [52, 53].
The input of this program is the scattering phase extracted from numerical data as denoted in Section 1: The scattering phase $\phi(k_{1},k_{2})$ vanishes for non-interacting particles, and approaches $\pi$ for hard-core bosons; for truly interacting cases, it has a non-trivial momentum dependence. Indeed, its momentum-derivative $\partial_{k_{1}}\phi(k_{1},k_{2})$ directly connects to the displacement observed after scattering [48].
The determination of the effective interactions $U_{\text{2-eff}}$ from the scattering phase $\phi(K,q)$ is the celebrated inverse scattering problem [51]: the solution is unique, and can be numerically found evaluating the Gel’fand-Levitan-Marchenko’s equations [52] or solving an optimization problem based on chosen trial potentials [53]. Instead of aiming for a full characterization of the effective interactions we can, however, focus on the emergent theory in the assumption that the momenta of the scattering magnons belong to a small shell $k\in I_{p}\equiv[p-\delta p,p+\delta p]$. In this approximation, we Taylor expand the dispersion law
$\varepsilon(k_{1,2})\simeq\varepsilon(p)+v(p)(k_{1,2}-p)+\frac{1}{2m^{*}(p)}(k_{1,2}-p)^{2}+\mathcal{O}\big{(}(k_{1,2}-p)^{3}\big{)},$ (S4)
and approximate the interaction $U_{2-\text{eff}}(k_{1},k_{2})\simeq U_{2-\text{eff}}(p,p)$. Within this approximation, it is easy to convert $H_{\text{2-eff}}$ into a real space Hamiltonian, where the spatial coordinates $x_{1,2}$ conjugated to the small momenta shift $k_{1,2}-p$ are approximated as continuous
$\Delta\mathcal{H}_{\text{2-eff}}\simeq-\tfrac{1}{2m^{*}(p)}(\partial^{2}_{x_{1}}+\partial^{2}_{x_{2}})+iv(p)(\partial_{x_{1}}+\partial_{x_{2}})+c(p)\delta(x_{1}-x_{2})\,,$ (S5)
where $\Delta\mathcal{H}_{\text{2-eff}}=\mathcal{H}_{\text{2-eff}}-2\varepsilon(p)$ and $c(p)\equiv U_{\text{2-eff}}(p,p)$. The Hamiltonian (S5) describes the two-particle sector of the Lieb-Liniger (LL) model [54, 55]: with a contact potential, in a moving frame with velocity $v(p)$. The scattering phase of this model is $\phi_{\mathrm{LL}}(k_{1},k_{2})=-2\arctan\left(\tfrac{k_{1}-k_{2}}{m^{*}(p)c(p)}\right)$ [54, 55]: by comparing the true scattering phase with the effective one, we can determine the effective interaction. Since we are within a small relative momentum approximation, we fix the effective interaction $c(p)$ through the scattering length $a(p)\equiv\lim_{k_{1}\to p}\partial_{k_{1}}\phi(k_{1},p)$, and the LL correspondence $a(p)=-1/(m^{*}(p)c(p))$.

### v2

@@ C.1 Scattering Phases from Two Magnon Spectra
C.2 Comparison of Real and Fourier Space Results
Figure 6: Scattering Phases and Shifts from Real Space Simulations. (a) We show scattering phases $\phi_{q}(\Delta)$ for two magnon scattering with total momentum $K=0$ and relative momenta $q\in\{\frac{\pi}{32},\frac{2\pi}{32},\frac{3\pi}{32}\}$ extracted from real space evolution. Results are set into comparison to XXZ predictions (solid lines). (b) Scattering phases $\phi_{q}(\Delta)$ for comparable parameters are extracted from the two-magnon spectra. The values of $q$ are thereby obtained from the fitting process and are approximately given by $q\in\{0.1,0.2,0.3\}$. (c) Scattering shifts $d_{1},d_{2}$ extracted from real space simulations of magnons pairs $(K,q)=(-\frac{\pi}{8},\frac{\pi}{16})$ agree well with analytical predictions (solid lines) within the finite spatial resolution of the lattice. Error bars indicate the standard deviation of measurements of $d_{1},d_{2}$ at different times after the scattering event. (d)-(e) We show analytical predictions for the relative scattering shift $\Delta d=d_{1}-d_{2}$ of two magnons following the XXZ description for anisotropies $\Delta=0.1$ (b) and $\Delta=0.8$ (c) in units of the lattice spacing $a$. We focus on the region of the Brillouin zone where chiral magnon bands are located, i.e. $k\in[-\frac{\pi}{4},\frac{\pi}{4}]$.
[…]
To furthermore connect to experiment, we can use the magnon profile of real space evolutions to directly determine the measured scattering shifts $d_{1},d_{2}$ of the first and second magnon, respectively. The extracted scattering shifts for a magnon pair of total momentum $K=0$ and relative momentum $q=\frac{3\pi}{32}$ are shown in Fig. 6 (c) and set into comparison to $d_{1}=\partial_{k_{1}}\phi_{\text{XXZ}}(k_{1},k_{2})$ and $d_{2}=\partial_{k_{2}}\phi_{\text{XXZ}}(k_{1},k_{2})$ containing the XXZ scattering phase of Eq. (C.1). Indeed, we find characteristics of the analytical predictions clearly resolved in the experimentally accessible $d_{1},d_{2}$. Interestingly, the scattering shifts $d_{1},d_{2}$ deviate more strongly from XXZ predictions than the phases $\phi_{q}(\Delta)$ of Fig. 6 (a). This fact that can be attributed to the fitting process used to determine $d_{1},d_{2}$. For this, local peaks of the real space magnon distribution, indicating the magnon position, are extracted from simulations at multiple times after the scattering event and set into comparison with the free evolution of given magnon velocity. The standard deviation for measurements at different times is indicated by the error bars shown in Fig. 6 (c). We find a finite spatial resolution of the real-space lattice, i.e., given by the lattice spacing, to be one crucial aspect limiting the accuracy of the extracted scattering.
Figure 7: Isotropic Magnon Interactions. (a) Scattering phases $\phi$ for magnon scattering with momenta $(K,q)=(-\frac{\pi}{8},\frac{\pi}{16})$ and interwire magnon interaction $\Delta_{y}\in\{0.05,0.1,0.15\},J_{y}=0.5$ as a function of intrawire interaction $\Delta$. For comparison we show predictions for the XXZ scattering phase $\phi_{\mathrm{XXZ}}$ (gray line). (b) Scattering shifts extracted from real-time simulations for the isotropic system $\Delta_{y}=1.0,J_{y}=1.0$. Predictions for XXZ are plotted for comparison (gray lines).
Appendix D Experimental Details
In our work, we propose an experimental protocol aiming to probe the scattering shift $\Delta d$ from interactions between two edge magnons using inelastic scanning tunneling microscopy (STM). In the following, we want to judge whether the current experimental setups allow for a sufficiently high resolution to measure the relative scattering shift $\Delta d=d_{1}-d_{2}$. Important contributions to this experimental resolution are (i) the width and (ii) the momentum spread of the magnon wavepackets injected in the beginning, and last but not least (iii) the spatial resolution $\Delta x$ of the detection. To estimate the magnitude of theses effects we perform valuations with parameters taken from current experimental platforms. As the width of the initialized wavepacket is expected to directly correlate with the typical range of interactions between tip and sample, which also determines the spatial resolution, we expect both effects (i) and (iii) to be on the same order of magnitude. For small sample-tip distances of $h$ ($\sim 0.5-1.5\text{nm}$) and sharp tips of radius $R$ ($\sim 0.8\text{nm}$) the resolution can be computed within the Tersoff-Hamann model as $\Delta x=[(0.2\penalty 10000\ \text{nm})(R+h)]^{\frac{1}{2}}\sim 0.5\text{nm}$ [63]. Next we have to analyze the momentum window $\Delta k$ excited due to finite energy resolution $\Delta E$ in STM. For the latter experiments working at high energy resolution demonstrated an accuracy of $\Delta E=0.1\text{meV}$ [64, 65, 66]. Assuming a more generous estimate of $\Delta E=0.25\text{meV}$ and typical velocities of chiral magnons of $v\sim 1000\frac{m}{s}$, e.g. for van-der-Waals magnets of lattice constant $a\sim 1\text{nm}$ [67, 31, 30], this relates to a momentum uncertainty of $\Delta k=\frac{\Delta E}{\hbar v}\sim 0.38\frac{1}{\text{nm}}$. In conclusion the initialized wavepacket will spread over $6\%$ of the Brillouin zone and potentially important broadening of the local wavepackets only appears if the dispersion shows non-negligible curvature over this range in momentum space. All effects combined hence cause an effect on the order of the lattice spacing $a$. What remains to be checked is if the scattering shift is expected to be resolvable taking into account the finite width of wavepackets and limited spatial resolution. For this we show analytical results for the XXZ scattering phase of Eq. (C.1) for different interaction parameters $\Delta\in\{0.1,0.8\}$ in Fig. 6 (d) and (e). We find that the observed scattering shift $\Delta d$ typically takes values of a few lattice spacings $a$ and hence is within the resolution of current platforms.
Another aspect certainly relevant for experiment is the tunability of the bare interaction parameter $\Delta$. While in principle the latter can be tuned by applying external strain to the sample, there is an alternative way to cause resonances between edge modes and bulk bound modes in the system. Instead of tuning $\Delta$ variation of the Aharonov-Casher phase $\delta\varphi$ can lead to similar phenomena as demonstrated in the main text. Changes of $\delta\varphi$ do not shift the bands of bulk bound magnons in energy, but instead translate the momentum window of the chiral magnon band $\varepsilon(k)$ and with that move the energy-momentum window for two chiral edge magnons, see marked in blue in Fig. 3 a), within the Brillouin zone.
Appendix E Isotropic Interactions
While in the main text we discuss the limiting case of magnon interactions taking place only within the individual wires a more realistic scenario might include interactions also across different wires. To demonstrate that our analysis also applies in similar fashion to this scenario we generalize the coupling between chains and consider the Hamiltonian
$\mathcal{H}=-\frac{J_{x}}{2}\sum_{x,y}\Bigg[\Big(e^{-i\varphi_{y}}\sigma_{x,y}^{+}\sigma_{x+1,y}^{-}+\mathrm{h.c.}\Big)+\Delta\sigma_{x,y}^{z}\sigma_{x+1,y}^{z}\Bigg]-\frac{J_{y}}{2}\sum_{x,y}\Bigg[\Big(\sigma_{x,y}^{+}\sigma_{x,y+1}^{-}+\mathrm{h.c.}\Big)+\Delta_{y}\sigma_{x,y}^{z}\sigma_{x,y+1}^{z}\Bigg].$ (11)
In contrast to the previously discussed case, the presence of finite interactions $\Delta_{y}$ is expected to change the interacting theory on the edge already in the perturbative limit of weakly coupled chains ($J_{y}\ll 1$). This, in particular makes a comparison to analytical $XXZ$ results inapplicable. Scattering properties can, however, be analyzed following the same strategy as before. We show results for scattering phases $\phi$ for weak interwire interactions $\Delta_{y}\in\{0.05,0.1,0.15\},J_{y}=0.5$ in Fig. 7 (a) and scattering shifts $d_{1},d_{2}$ for isotropic coupling $J_{y}=1.0,\Delta_{y}=1.0$ in Fig. 7 (b), respectively. Following the remarks of Section IV one can determine the effective interation strength at the edge. Relating the observed scattering properties to resonances with bulk modes would, however, require a separate analysis due to changes in the nature of bulk bound states due to finite interactions $\Delta_{y}$.